OpenStudy (anonymous):

Given the linear equation 3y + 4x =12, find the y-intercept. A. 4 B. 09_01_ST_12_aa.gif C. –4 D. 09_01_ST_12_ac.gif

OpenStudy (anonymous):

B. -4/3 D. 4/3 @zeinas.

OpenStudy (anonymous):

@wolf1728

OpenStudy (anonymous):

Hi. My computer timed out. So first, we need to subtract 4x from both sides: 3y+4x-4x=-4x+12. Cancel out the the both 4x on the left side. What do you get?

OpenStudy (haseeb96):

y=mx+b for y inter cept put x=0 3y=-4x+12 3y=12 y=4

OpenStudy (haseeb96):

A.4 is the correct answer

OpenStudy (anonymous):

3y=-4x+12 y=-4/3x+4 so the y intercept is b in y=mx+b so its 4

OpenStudy (anonymous):

ok thanks just one more question and I would be forever gratefull!!!!

OpenStudy (anonymous):

Solve for x. 09_01_ST_14_b.gif A. x = 2 B. x = 41 C. x = 6 D. x = 43

OpenStudy (anonymous):

Sure. Are you going to post it here or tag me in?

OpenStudy (anonymous):

x-1/12=7/2

OpenStudy (anonymous):

@zeinas.

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

You are the best teacher ever!!! I hope you know how thankful I am!!!! @zeinas.

OpenStudy (anonymous):

Thanks.:) So to solve that we need to add 1/12 to both sides of the equation. so it will be x+1/12-1/12=7/2 +1/12 What will we have when 1/12 cancel from the left side?

OpenStudy (anonymous):

x=7/2+1/2?

OpenStudy (anonymous):

x=7/2+1/2

OpenStudy (anonymous):

I think its 6?

OpenStudy (anonymous):

That will be right. But I do have a question. Was your question x-1/12=7/2 or x-1/2=7/2?